INAF LBC+MODS queue observing and AO PISCES, 2012 June 20-21

Observer: C. Arcidiacono, R. Briguglio, K. Boutsia, M. Faccini, F. Vitali
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC + PISCES

Summary

Clear skies at the beninng of the night and the average wind is 3m/s. We had <1.0" seeing for the whole night

The night starts with LBC and dimm seeing 0.9".
Right from the statrt we had problems collimating on the blue channel because we would reach limits for the mirror on the blue side.
40min lost because of problems with the pointing procedure.
Slewing to the science field we kept having poor collimation in the blue and although stars were round in the central chip, they were quite elongated on the other chips even right after dofpia.
The red tracker stopped during the second OB and we lost 30min to run LBCkill/LBCstart successfully.
The second time the red stopped, we had already reached the limit in airmass for Nonino, and we decided to just switch to Pisces, without trying to recover.

At 07:35UT we switch to pisces.
The seeing was very good today and we obtained some extra data for the Cesaroni target.
20min lost due to an AGW error - we had to power cycle AGW and restart the AZcam server.
We spent the second half of the night, obtaining ADI data for Mesa (Hip101800) and the program is finished.
At the end of the night the ND was installed to get some unsaturated frames of the target.
The AO system today was stable.
At the end of the night we got a good set of FeII flats.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Nonino_19 EGS Us r 1h30min seeing measured on central chip blue:1.0"/red:078" - allseeing for the whole mosaic blue:1.2"/red:0.86"
Standard SA104 Us/Bbess - rsloan 10min airmass 1.3
Standard SA109 Us/Bbess - rsloan 10min airmass 1.7
Cesaroni IRAS20126 Ks, H2, Br-g 5min for Ks / ~15min for Brg/H2 since tonight the seeing is excellent we obtained some extra data of better quality
Mesa HIP101800 H, FeII 500frames in 1hr / 2min in FeII+ND ADI data + data with ND + calibration star - COMPLETED

Details

01:24 TurnON LBC - connect to LBT - 10Bias_Bino

02:50 start Nonino 19B

02:53 flat Us and r

03:09 dofpia, /first/, /x2 on ACT0188 - Saturated

03:13 dofpia, /first/, on ACT0188 - Signal still high

03:18 dofpia, /first/, on ACT0188 - Still high

03:23 dofpia, /first/, on ACT0188 - Still high

03:30 dofpia, /first/, on ACT0188

03:38 lbcrangebal

03:42 Standard Sa 104 (airmass 1.318)
  • Blue images are elongated - collimation problem - PT

03:49 Standard Sa 109 (airmass 1.72)
  • dofpia ./x2

04:00 Standard Sa 109 (airmass 1.72)

04:08 Focus
  • dofpia ./x2 didn't converge
  • dofpia ./x2 converge

04:20 EGS focus
  • lbcrangebal (bad command to align)
  • sent ofset for wrong images number
  • reset telescope to restart .....
  • dofpia on target to reduce alignement

04:41 Send EGS focus again
  • star not in field
  • need to restart the telescope again

04:50 Send again EGS focus
  • dofpia ,/x2
  • lbcrangebal with right images number
  • collimation and dofpia are good

05:02 send nonino EGS Us r 3

05:41 dofpia

05:45 EGS Us r 4

06:05 The red tracker has stopped and the OB is hung ...
  • we had to run LBCkill/LBCstart

06:22 diofpia ,/x2

06:27 EGS Us r 4 - again
  • the is elongation in blue - more obvious on the fourth chip
  • alert messages
3	2012/06/21	07:02:10.995297	TRACKERS		RPC Client/Host communication failed retval:5 [src/trackers/trackers.c:1625]
2	2012/06/21	06:59:41.248888	TRACKERS		RPC Client/Host communication failed retval:5 [src/trackers/trackers.c:1625]
1	2012/06/21	06:57:10.374615	TRACKERS		RPC Client/Host communication failed retval:5 [src/trackers/trackers.c:1625]
0	2012/06/21	06:54:40.547500	TRACKERS		RPC Client/Host communication failed retval:5 [src/trackers/trackers.c:1625] 

07:07 dofpia ,/x2

07:10 EGS Us r 5

07:32 The red tracker stopped AGAIN!!!! We need to run LBCkill/LBC start - the source is at the limit for the airmass specification
  • we decide to switch to pisces

07:35 SWITCH TO PISCES

08:07 Send the first preset - IRAS20126 (Cesaroni)

08:15 loop closed - start obtaining data

09:05 problems with the guider - before going to the next target the TO power cyclse AGW
  • no real time lost, since we go on taking sky frames in the meantime

09:14 still problems with the guider - restarting the AZcam server

09:26 we send an active preset to see if the problem has been fixed
  • success

09:30 we send the preset again - this time ACE
  • we cannot find the star on the pyramid
  • Steve has to center the star manually after we send an acquire preset

09:36 Send ACE preset to HIP101800 (Mesa)
  • loop closed

09:42 start obtaining science data

10:41 open the loop manually and stop the script
  • send the telescope to zenith in order to install the ND
  • obtain sky frames in the meantime

10:54 send preset again to get the data with the ND
  • send an acquire preset so that Steve can manually center the guide star

10:57 send the ADI preset
  • we try different exptimes and filters to see what combination gives not saturated images

11:10 run autofocus since we have the ND on
  • we got the focused images that we needed

11:24 slew to AO699 for calibration

11:34 loop closed - getting calibration data +sky

11:40 slew to get sky frames

12:00 try to get some FeII flats - counts from 500ADU to 9000ADU

12:47 script for darks

-- KonstantinaBoutsia - 21 Jun 2012
Topic revision: r5 - 22 Jun 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback